OpenStudy (yanasidlinskiy):

The width w of a rectangular swimming pool is 7x^2. The area A of the pool is 7x^3 – 42x^2. What is an expression for the length of the pool? This is what I have done so far. W=7x^2 7x^3-42x^2=7x^2(x-6)

OpenStudy (yanasidlinskiy):

@AccessDenied @beccaboo333 @whpalmer4 Help!!!

OpenStudy (whpalmer4):

\[A = lw\]right?

OpenStudy (yanasidlinskiy):

Yes..

OpenStudy (yanasidlinskiy):

I don't know what to do next after that equation part..Like solving it.

OpenStudy (whpalmer4):

if you know how to do polynomial division, you could find \[l = \frac{A}{w}\] or if you've factored \(A\) and one of the factors is \(W\), then the other factor must be \(l\)

OpenStudy (yanasidlinskiy):

I have no clue how to do that.

OpenStudy (whpalmer4):

Okay, well you started to factor A, completed it, even. \[A = 7x^2(x-6)\]If \(A = l*w\) and \(w = 7x^2\), doesn't that mean \[A = w(x-6)\]?

OpenStudy (yanasidlinskiy):

Yes...I guess..idk..

OpenStudy (yanasidlinskiy):

Would it be \[7x^2(x-6) l* 7x^2\]

OpenStudy (whpalmer4):

It does :-) \[A = 7x^3 – 42x^2= 7x^2(x-6)\] \[A = l*w\]\[w = 7x^2\] \[A = 7x^2*l\]\[7x^2(x-6) = 7x^2*l\]\[\frac{7x^2(x-6)}{7x^2} = \frac{7x^2*l}{7x^2}\]\[\frac{\cancel{7x^2}(x-6)}{\cancel{7x^2}} = \frac{\cancel{7x^2}*l}{\cancel{7x^2}}\] \[l=(x-6)\]

OpenStudy (yanasidlinskiy):

Ok. I see what you did! Makes sense now:) Thank you so much!!!!!

OpenStudy (whpalmer4):

You're welcome!

OpenStudy (whpalmer4):

It seems so obvious to me, I want it to be that way for you, too :-)

OpenStudy (yanasidlinskiy):

Yea..Like I need it step by step...At least somebody helped me solve this!